OpenStudy (anonymous):

A car going 70 mph stops in 145 feet. Find the acceleration, assuming it is constant. Give your answer in feet per second-squared.

OpenStudy (osanseviero):

You should post this in the physics group.

OpenStudy (anonymous):

Yet, this is from my math class. This teacher gives me brain cramps.

OpenStudy (amistre64):

acceleration is defined as change in velocity over time so (vf - vi)/time

OpenStudy (amistre64):

hmm v = dt t = d/v; ft/(ft/sec) = sec so the dimensions work fine

OpenStudy (amistre64):

gonna have to convert miles/hr to feet/sec prolly .. or vice versa

OpenStudy (amistre64):

get my mind right velocity is a measure of miles /hour v = d/t therefore t = d/v had it right in my head

OpenStudy (amistre64):

we can work it backwards too, start at 0 and speed up to 70mph over 145 feet .... a = k v = kt + c , v(0) = 0, c=0 kt = 70mph, when t=70mph/k d = 1/2 kt^2 + j, d(0) = 0 so j=0 d(70mph/k) = 145feet 145ft = 1/2 k(70mph/k)^2 290ft = (70mph)^2/k k = (70mph)^2/290ft

OpenStudy (amistre64):

convert as needed :)